The following data are from a completely randomized design. Treatment 163 145 123 142 157 121 166 129 132 144 145 144 148 138 153 191 138 131 159 142 134 344.8 88.8 152.8 Sample mean Sample variance a. Compute the sum of squares between treatments. Round the intermediate calculations to whole number. b. Compute the mean square between treatments. c. Compute the sum of squares due to error. d. Compute the mean square due to error (to 1 decimal).
Answer #1

The formula of sum of square between treatments is as follows:

SS_{b}=sum_{i=1}^{k}n_{i}*(ar X_{i}- ar {ar X})^{2}

Therefore,

X (1/3) * (1594 142+ 134) 145

SSb = 6*(159-145)^2 + 6*(142 - 145)^2 + 6*(134 - 145)^2 = 1176 + 54 +726 = 1956

Mean square between treatments = SSb / (k - 1) = 1956/(3-1) =1956/2 =978

d) Let's find mean square due to error

MSE =-* (s[+ s s )-* (344.84 88.84 152.8) = 195.4667

After rounding we get MSE = 195.5 ( this is the answer of part s

C) Sum of square due to error ( SSE)

SSE = ( n - k ) * MSE = ( 18 - 3) * 195.4667 = 2935 ( This is the answer of part c).